# Scaling-Up the Pretraining of the Earth Observation Foundation Model PhilEO to the MajorTOM Dataset

### v1

@@ IV Related Work
Recent studies on GFMs, summarized in Table I, highlight the increasing potential of learning robust feature representations from large-scale Earth Observation (EO) data, later used for fine-tuning on various downstream tasks. However, most existing studies have been limited by relatively modest dataset sizes and have not explored training GFMs at the scale of more than 20TB. This motivates our work, where we develop and pretrain the PhilEO MajorTOM model on the full 23TB MajorTOM Core-S2L2A dataset. In addition, few-shot learning has gained attention within the EO community, as it enables models to generalize effectively from limited labeled data. Prior work [17, 18] demonstrated the advantages of few-shot learning for land cover classification, showing that models can maintain strong performance even with a small number of examples.
The synergy between Foundation Models and few-shot learning was further explored in [19], showcasing improved segmentation performance through combined training strategies. Additionally, frameworks such as [20] have proposed systematic approaches to low- and few-shot adaptation in the context of Foundation Models.
@@
In addition, to this day, U-Net models are still considered the default benchmark for semantic segmentation, often outperforming new architectures. [16] benchmarked the performance of Swin-UPerNet against a U-Net on semantic segmentation use cases. It was shown that the Swin-UPerNet model is a good competitor for the U-Net.
@@
V Evaluation and Results
We evaluate the proposed models, comparing different architectures and analyzing the impact of scaling-up pretraining. Our evaluation focuses on the three PhilEO Bench downstream tasks: road density estimation, building density regression, and land cover mapping. Before training the models directly on MajorTOM, we first train them on the 0.5TB PhilEO Global dataset in order to select the best architecture to scale-up.
V-A PhilEO ViT UPerNet pretrained on PhilEO Globe 0.5TB
As shown in Fig. 2, the PhilEO ViT UPerNet model consistently outperforms the PhilEO ViT model with a convolutional decoder across all $n$-shot settings for road density estimation. Notably, the PhilEO ViT UPerNet achieves comparable performance to the PhilEO Geo-Aware U-Net model [3], highlighting the effectiveness of the UPerNet decoder for pixel-wise regression tasks.
In Fig. 3, the PhilEO ViT UPerNet demonstrates improved RMSE performance for building density estimation compared to the convolutional decoder ViT baseline. For instance, in Fig. 3, at $n=50$, the PhilEO ViT UPerNet achieves a RMSE of $0.08944$, compared to $0.2$ for the convolutional decoder ViT model, corresponding to an improvement of $55.28$%. At $n=100$, the ViT UPerNet model further improves to a RMSE of $0.08367$, versus $0.1$ for the convolutional decoder ViT baseline, leading to a $16.33$% relative gain.
@@ V Evaluation and Results
In Fig. 4, the PhilEO ViT UPerNet also outperforms the ViT CNN model across all $n$-shot experiments for land cover mapping, in accuracy. Moreover, both the PhilEO ViT UPerNet and the Geo-Aware U-Net outperform a fully supervised U-Net model (without pretraining), particularly in low-shot regimes. Overall, the PhilEO Geo-Aware U-Net achieves slightly better performance on average than the PhilEO ViT UPerNet across the three downstream tasks, as shown in Figs. 2-4.
V-B Scaling-up: Evaluation of PhilEO MajorTOM 23TB GFM
@@
According to the results in Section V-A, we scale-up the best-performing architecture, the PhilEO Geo-Aware U-Net, and we pretrain it on the full 23TB MajorTOM dataset, as well as on the FastTOM 2TB specialized subset. We study the scaling-up to 23TB (and 2TB) data, as well as the number of parameters 200M (versus 44M), and the architecture, i.e. ViT UPerNet (versus Geo-Aware U-Net [3]).
@@
Road density estimation. Fig. 5 shows the RMSE performance for road density estimation across different $n$-shot settings. For all $n$-shots, the scaled-up PhilEO MajorTOM 23TB model outperforms the Geo-Aware U-Net model pretrained on the PhilEO Globe 0.5TB dataset. At $n=100$, the PhilEO MajorTOM 23TB model achieves a RMSE of $0.06724$, compared to $0.07070$ for the model pretrained on the PhilEO Globe 0.5TB dataset. This corresponds to an improvement of $4.89$%. For $n=50$ and $100$, the PhilEO 44M MajorTOM model outperforms the FastTOM and PhilEO Globe pretrained baselines, demonstrating the benefits of large-scale pretraining even when fine-tuning with limited labeled data.
@@
Both larger models show strong performance, with PhilEO 200M FastTOM outperforming the others at several $n$-shots.
@@
Building density regression. As shown in Fig. 6, similar trends are observed for building density estimation. At $n=100$ and $500$, both the PhilEO 44M MajorTOM and FastTOM models outperform in RMSE the PhilEO Globe 44M baseline. In addition, the PhilEO 200M FastTOM model again performs best across most $n$-shot settings.
@@
To better understand the impact of specialized versus general-purpose pretraining, we compare the performance of models trained on specialized land-only datasets (PhilEO Globe 0.5TB and FastTOM 2TB) with the performance of models trained on a large general-purpose dataset (MajorTOM 23TB) which includes oceans and ice. Given that all models share the same Geo-Aware U-Net architecture (approximately 44M parameters) and that the smaller datasets use a land mask to exclude oceans and ice, we expect models trained on PhilEO Globe and FastTOM to perform better in downstream land-focused tasks. Indeed, in most cases, this expectation holds. However, for the road density estimation task under low-shot (few-shot) learning settings, the model pretrained on MajorTOM surprisingly outperforms the specialized models in terms of RMSE. This suggests that despite the broader and noisier pretraining data, large-scale general pretraining can be advantageous under certain conditions.
This study allows us to analyze two key factors:
  • Specialization effect: Does pretraining on land-only images give an advantage when downstream tasks are also land-focused?

### v2

@@ III-C Mamba 2D Models
In contrast to Mamba SSM models, instead of the selective scan mechanism, Transformers (including ViT) use attention. For Query, Key, Value (Q, K, V) multi-head attention [30]:
$\textbf{S}_{Q}=\textbf{x}\textbf{W}_{Q}^{T}\text{,}\ \textbf{S}_{K}=\textbf{x}\textbf{W}_{K}^{T}\text{,}\ \textbf{S}_{V}=\textbf{x}\textbf{W}_{V}^{T}$ (9)
where we observe that we have trainable linear projections (like, for example, in (5)-(8) for the input-dependent matrices $\textbf{S}_{\textbf{B}}$, $\textbf{S}_{\textbf{C}}$, and $\textbf{S}_{\Delta}$). In (9), $\textbf{W}_{Q}$, $\textbf{W}_{K}$, and $\textbf{W}_{V}$ have dimensions $\mathbb{R}^{N\times D}$. Next, for ViT, the attention weights are computed; these are given by: $S=\text{Softmax}(\textbf{S}_{Q}\textbf{S}_{K}^{T}/\sqrt{d_{k}})\,\textbf{S}_{V}$. Here, the scaled dot product uses the dimension of the Key vectors, $d_{k}$. Multi-head attention [9] involves also using concatenation of the $m$ attention weights and a final learnable linear transformation to project the multi-head attention scores to the outputs.
ViT models have significantly less image-inductive bias compared to CNN U-Net models. This is because ViT attention correlates everything with everything else. Moreover, in self-attention, image token order does not matter [9], [30]. In addition, Mamba models have more image-inductive bias compared to ViT models, as in Mamba S6 SSM models, the image token order is important and matters.
III-D The S12-PhilEOBench evaluation framework
In this work, we also extend the PhilEO Bench dataset [3] to include co-located S-1 SAR data in addition to S-2 multi-spectral images. In Figs. 2 and 3, we show example images [^3] The S-1 SAR PhilEO Bench data can be found in http://huggingface.co/datasets/NikolaosDionelis2023/s1-phileobench/tree/main. In addition, we note that the S-2 multi-spectral PhilEO Bench data can be found in http://huggingface.co/datasets/NikolaosDionelis2023/s2-phileobench/tree/main. from the global dataset S12-PhilEOBench (for example, Fig. 2 is from East Africa). For the same S-2 and S-1 images, we have three types of labels: road density pixel-wise regression (continuous real-valued output), building density regression, and land cover pixel-wise classification (with $11$ classes).
@@ IV Related Work
Recent studies on GFMs, summarized in Table I, highlight the increasing potential of learning robust feature representations from large-scale EO data, later used for fine-tuning on various downstream tasks. For a recent study on the scope of geo-spatial FMs, we refer the reader to [60]. However, most (but certainly not all) existing studies have been limited by relatively modest dataset sizes and have not explored training GFMs at the scale of more than 20TB, and this is important. This motivates our work, where we develop and pretrain the PhilEO MajorTOM model on the full 23TB MajorTOM Core-S2L2A dataset. Furthermore, during the ESA-NASA Workshop on AI Foundation Models for Earth Observation [61] (in May 2025), it was highlighted that the performance of FMs is not uniform across tasks; moreover, in certain cases, conventional models surpass state-of-the-art EO FMs. This observation prompts a critical inquiry: should the field pursue a single universal model, develop application-specific models, or adopt a Mixture of Experts (MoE) strategy?
The recent introduction of Meta’s improved and upgraded DINOv3 framework [62] provides an additional rationale for enlarging the dataset scale. Although the large-scale proprietary geo-spatial dataset SAT-493M—comprising $493$ million images at 0.60.6m resolution—has not been publicly released, its storage footprint can be reasonably estimated to substantially exceed 2020TB, and this is important. It has been demonstrated that the $7B$-parameter variant of the DINOv3 model, pretrained on the SAT-493M dataset, attains state-of-the-art performance across $12$ of $15$ benchmark tasks, including classification, segmentation, and horizontal object detection. Evaluated on the Geo-Bench suite, DINOv3 surpasses prior leading approaches such as Prithvi-EO-2.0. Furthermore, under a linear probing paradigm—where the backbone (EO FM) is frozen—DINOv3 establishes superior results on three previously unsaturated classification tasks and on five out of six segmentation tasks, further underscoring its robustness and generalization capacity. This indicates that, for specific important downstream applications, general-purpose self-supervised learning can serve as a competitive alternative to domain-specialized fine-tuning in the context of geo-spatial data.
The main shortcomings of existing methods, also in continuation of the discussion in Secs. II and I, as well as in continuation of Table I, is the quadratic complexity of ViT and attention with respect to the image token sequence length [34], [35], the potential performance benefits of the Mamba S6 2D SSM model (i.e. see Sec. V), seeing how much faster is for EO data Mamba compared to ViT, the effect of scaling-up the pretraining dataset size to 2323TB (as well as to 22TB), and the impact of increasing the model size (i.e. number of parameters) on the model’s final performance on the downstream tasks.
TABLE II: The main models we develop, implement, train, and evaluate in this work. We present and categorize the models in this Table. The models are listed in order of appearance in the figures.
Name of models, including sizes Figures in Subsection(s) in
ViT with CNN decoder, 300M-0.5T 4-6 V-A
ViT with UPerNet decoder, 300M-0.5T, 100M-2T, 100M-23T 4-11, 15-17 V-A-V-C, V-E
Geo-Aware U-Net, 44M-0.5T, 200M-2T, 44M-23T, 44M-2T 4-11 V-A-V-C
RS3Mamba 168M-2T 12-14 V-D
Mamba UPerNet P16 100M 12-14 V-D
TerraMind, v1.0-B, v1.0-L [29] 18-20 V-E
Prithvi-EO-2.0 600M TL [23] 18-20 V-E
@@
We note that because many EO FMs have been proposed recently, about the choice of models in the summary Table I, for the model proposed in [4], the code is not available. In addition, to this day, U-Net models are still considered the default benchmark for semantic segmentation, often outperforming new architectures. Moreover, [16] benchmarked the performance of Swin-UPerNet against a U-Net on semantic segmentation use cases. It was shown that the Swin-UPerNet model is a good competitor for the U-Net.
Furthermore, few-shot learning has gained attention within the EO community, as it enables models to generalize effectively from limited labeled data. Prior work [17, 18] demonstrated the advantages of few-shot learning for land cover classification, showing that models can maintain strong performance even with a small number of examples. The synergy between Foundation Models and few-shot learning was further explored in [19], showcasing improved segmentation performance through combined training strategies. Additionally, frameworks such as [20] have proposed systematic approaches to low- and few-shot adaptation in the context of Foundation Models.
V Evaluation and Results
In this section, we present the evaluation and results of the models when we scale-up EO FMs in terms of larger pretraining dataset, i.e. the MajorTOM Core-S2L2A 23TB dataset, larger model, and architecture. According to the results, evaluation, and experiments, EO FMs pretrained on general-purpose data, when scaled-up appropriately, can eventually outperform models pretrained on specialized data. We are interested in Earth Observation, and not only in Land Observation or Ocean Observation (or Ice Observation). In the literature, there exist FMs only for oceans or ice, for example HydroFM [25]. When we pretrain on the entire MajorTOM 23TB dataset, we do not use any prior information, while when we pretrain on specialized datasets that have only land and exclude oceans, ice, and deserts, i.e. the FastTOM 2TB dataset, then we use prior information. The downstream tasks we are interested in are on land, i.e. PhilEO Bench [3]: road density estimation, building density regression, and land cover mapping [^4] The PhilEO Bench downstream tasks labeled dataset (0.40.4TB) was presented in [3], and we also note that a small and more manageable subset can be found in http://huggingface.co/datasets/JenteBosmans/tiny_phileo.. We focus on land downstream tasks and not on oceans or ice (or deserts), and this is why specialized datasets have prior knowledge.
@@ V Evaluation and Results
In this paper, we have developed and implemented several models, including for example the PhilEO Geo-Aware U-Net 200200M model, as well as the ViT UPerNet 100100M-2T model, the RS3Mamba 168168M-2T (e.g., in Fig. 14), and the patches-based Mamba UPerNet model which we will examine and evaluate in Sec. V-D. We list and summarize these main models of this paper in Table II, where we also summarize the main subsection in which the models are evaluated in. In this section, we evaluate the proposed models, comparing different overall model sizes and architectures (e.g., ViT vs. Mamba), analyzing the impact of scaling-up pretraining (i.e. Sec. III-A). The results are presented, examined, and discussed in the next subsections, i.e. in Secs. V-A-V-F.
Before training the models directly on MajorTOM, we first train them on the 0.5TB PhilEO Globe dataset [3] in order to select the best architecture to scale-up. Here, the main reason for this, is that it is expensive, requiring a lot of compute power, GPU machines, and GPU hours (i.e. real-world and pragmatic realistic scenario), as well as it is time-consuming, to train the models on the 23TB MajorTOM dataset. For these reasons, in the next subsection, we first present and examine the results of the models that we developed, implemented, and trained on the 0.5TB PhilEO Globe dataset.
V-A PhilEO ViT UPerNet pretrained on PhilEO Globe 0.5TB
As shown in Fig. 4, the PhilEO ViT UPerNet model consistently outperforms the PhilEO ViT model with a convolutional decoder across all the $n$-shot settings for road density estimation. Notably, the PhilEO ViT UPerNet achieves comparable performance to the PhilEO Geo-Aware U-Net model [3], highlighting the effectiveness of the UPerNet decoder for pixel-wise regression tasks. We test the models in the figures using the $n$-shot evaluation protocol, where we note that the horizontal x-axis is the $n$-shot number of samples per country, i.e. $6$ countries, $50$ times $6$ (i.e. $300$) data samples for the first point in the horizontal axis.
In Fig. 5, the PhilEO ViT UPerNet demonstrates improved RMSE performance for building density estimation compared to the convolutional decoder ViT baseline. For instance, in Fig. 5, at $n=50$, the PhilEO ViT UPerNet achieves a RMSE of $0.08944$, compared to $0.2$ for the convolutional decoder ViT model, corresponding to an improvement of 55.2855.28%. At $n=100$, the ViT UPerNet model further improves to a RMSE of $0.08367$, versus $0.1$ for the convolutional decoder ViT baseline, leading to a 16.3316.33% relative gain.
@@
In Fig. 6, the PhilEO ViT UPerNet also outperforms the ViT CNN model across all $n$-shot experiments for land cover mapping, in accuracy. Moreover, both the PhilEO ViT UPerNet and the Geo-Aware U-Net outperform a fully-supervised U-Net model (without pretraining), particularly in low-shot regimes. Overall, the PhilEO Geo-Aware U-Net achieves slightly better performance, on average, than the PhilEO ViT UPerNet across the three downstream tasks, as shown in Figs. 4-6.
V-B Scaling-up: Evaluation of PhilEO MajorTOM 23TB GFM
For clarity purposes, we plot and examine the performance of the models and our evaluation results in separate figures for: i) the Geo-Aware U-Net models (see Sec. III-A) and the ViT models (i.e. Sec. III-B), and also for ii) the best ViT model and the Mamba models. We focus on comparing no more than $5$ or $6$ models per plot, for clearness and clarity. By separating the plots in this way, we compare CNN-based U-Net models with ViT, which is an interesting comparison, and we also compare ViT with recent 2D Mamba S6 SSM models (see Sec. III-C), which is also a very interesting comparison and engenders cogent discussions mainly around the trade-off between computational complexity (i.e. ViT, attention, and quadratic complexity vs. Mamba, SSM selective scan [32], and linear complexity) and performance (i.e. lower RMSE or higher accuracy) in the final downstream task, for example for global S-2 road density estimation pixel-wise regression. In addition, U-Net CNN models also achieve linear complexity.
Regarding (i) above, the comparison of the models that are based on Geo-Aware U-Net and those that use ViT is for example in Fig. 7. In the following subsections, we will present both (i) and (ii), for example the results in Figs. 7-11 for (i), for the Geo-Aware U-Net models and the ViT models. In addition, we will also present and examine the results for (ii) above, for the best ViT model and the Mamba models (i.e. Sec. III-C), for example the results in Fig. 13.
@@ V-B Scaling-up: Evaluation of PhilEO MajorTOM 23TB GFM
In addition to the performance of the models on downstream tasks (e.g., in terms of lower RMSE and higher accuracy, for example Figs. 7-11), we are also interested in how much time do the models need to train. This is important, as we are interested in how much compute we use (i.e. GPU power and hours), and for example, to address the problem of quadratic complexity of the ViT (i.e. self-attention), many different methods have been proposed in the recent literature [42], including for example FlashAttention [43], linear attention [45], [46], partial attention [42], Swin Transformer [12], [9], MetaFormer which uses spatial pooling operations instead of attention [44], Mamba [36], [38], the Fast Fourier Transform (FFT) spectral block [47], [50], [59], and using attention at later stages and not at the initial model blocks [36].
For the evaluation of the models, continuing the discussion in the preceding paragraph, we count the FLOPs that are needed by the different models. The lower the FLOPs, the faster the model, i.e. the better (less operations). At this point, here, we also note that FLOPs are different from Floating-point operations Per Second (FPS). To be machine GPU agnostic, we use FLOPs and we do not measure time (i.e. seconds). In our FLOPs profiling analysis, we also include the models TerraMind [29] and Prithvi-EO-2.0 [23] (see Sec. V-E).
Sentinel-2, L2A data. For the models we develop, implement, train, and evaluate that are based on the EO FM PhilEO [3], we use images of size 128128x128128, and also $10$ spectral bands (channels), i.e. multi-spectral satellite Sentinel-2 data. More specifically, as in [3], we use the $10$ and $20$ metre spatial resolution spectral bands from the satellite Sentinel-2555http://sentiwiki.copernicus.eu/web/s2-mission#S2Mission-SpatialResolutionS2-Mission-Spatial-Resolutiontrue.
Image patches, and working with patches rather than pixels. In this paper, we also develop, train, and evaluate a Mamba S6 2D SSM model that is based on image patches, learnable linear projection, tokenization, and positional embeddings like in ViT. We use patches of size 1616x1616 (and we also experiment with the patch sizes of 88x88 and 44x44).
@@
Experiments and motivation. In this evaluation section, we present and examine the results of many models on the PhilEO Bench. The two recent models TerraMind [29] and Prithvi-EO-2.0 [23] are also evaluated on the PhilEO Bench in Sec. V-E. According to the results in Section V-A, we scale-up the best-performing architecture, the PhilEO Geo-Aware U-Net, and we pretrain it on the full 23TB MajorTOM dataset, as well as on the FastTOM 2TB specialized subset. We study the scaling-up to 23TB (and 2TB) data, as well as the number of parameters 200M (versus 44M), and the architecture—ViT UPerNet, i.e. see Sec. III-B— (vs. Geo-Aware U-Net [3]).
We study scaling-up and the behaviour and performance of the different models, and this is important. When we increase the size of the pretraining S-2 dataset to 23TB (i.e. Sec. III-A), we add ocean and ice data samples, and we effectively increase the aleatoric uncertainty, where here we note that aleatoric uncertainty is statistical and related to randomness, and the sample not being a typical example of the class and quantity we observe. Here, at this point, we also note that, in contrast to aleatoric uncertainty, epistemic uncertainty is systematic, caused by lack of knowledge (e.g., induced by the lack of detail in the measurements, for example by the lack of resolution or spectral information), and can be reduced by learning the specific characteristics of the quantity using additional information (for example, in-situ measurements).
Figure 8: Road and building density estimation: Evaluation in RMSE and also average over the different $n$-shots, examining the FLOPs needed by the models, where the best performing model is the PhilEO Geo-Aware U-Net. We also note that the two EO FM models TerraMind [29] and Prithvi-EO-2.0 [23] are also evaluated in Sec. V-E for roads, buildings, and land cover.
Figure 9: Land cover mapping: Evaluation in accuracy and average over the $n$-shots, examining the FLOPs of the models, where the best model is TerraMind.
V-C Comparison of scaling-up for U-Net and ViT models
@@
Road density estimation. Fig. 7 shows the RMSE performance for road density estimation across different $n$-shot settings. For all $n$-shots, the scaled-up PhilEO MajorTOM 23TB model outperforms the Geo-Aware U-Net model pretrained on the PhilEO Globe 0.5TB dataset. At $n=100$, the PhilEO MajorTOM 23TB model achieves a RMSE of $0.06724$, compared to $0.07070$ for the model pretrained on the PhilEO Globe 0.5TB dataset. This corresponds to an improvement of $4.89\%$. In the figure, we denote the Geo-Aware U-Net by U-Net for clarity purposes, where the Geo-Aware U-Net that we train and use is a modified U-Net pretrained on S-2 global data. For $n=50$ and $100$, the PhilEO 44M MajorTOM model outperforms the FastTOM and PhilEO Globe pretrained baselines, demonstrating the benefits of large-scale pretraining even when fine-tuning with limited labeled data.
@@
Both larger models show strong performance, with PhilEO 200M FastTOM outperforming the others at several $n$-shots. In particular, for $n$-shot $500$, for roads in Fig. 7, the Geo-Aware U-Net 200M (i.e. U-Net 200M-2T) has better performance than U-Net 44M (U-Net 44M-2T), and the percentage improvement is $4.31\%$, i.e. $0.061$ compared to $0.064$. Here, at this point, we also note that because jointly increasing the pretraining dataset size and the model size (i.e. scaling-up both the volume of the pretraining data (2323TB [15]) and also scaling the number of parameters) is important and also because pretraining on the entire MajorTOM dataset 2323TB is expensive, we defer as future work the training of the Geo-Aware U-Net 500500M 2323TB model (i.e. it is ongoing work).
@@
In Figs. 7-11, according to our FLOPs profiling analysis, the model Geo-Aware U-Net that is of size 4444M parameters has 5.875.87G FLOPs. This is low computational complexity (for example, see TerraMind and Prithvi-EO-2.0 in Sec. V-E). In addition, the Geo-Aware U-Net 200M has 82.2382.23G FLOPs. In Fig. 8, we compute, present, and examine the FLOPs needed by the different models, where the average over the two pixel-wise regression downstream tasks is calculated, as well as the mean over the different $n$-shots. The results show that, when taking into account the FLOPs (i.e. in Giga (G) FLOPs) that are required for the downstream tasks, the Geo-Aware U-Net model 200M (and 44M) have top/ very good performance.
The model Geo-Aware U-Net 44M has a small number of FLOPs, approximately 66G (Fig. 8), and this is one of the key takeaways of this paper. We observe that the small model wins in this case. About scaling laws for the models, we are interested in the improvement in performance when we increase the number of FLOPs. For example, Geo-Aware U-Net 4444M to 200200M in Fig. 8: $5.87$ to $82.23$ G FLOPs, RMSE $0.0698$ to $0.0653$. A relevant question that is important is which is the best model for a given cost, and whether saturation is observed in performance accuracy when increasing the number of FLOPs. In Fig. 8, FLOPs are for the downstream tasks (i.e. for the forward operation of the network for a single image), and for Mamba, see Sec. V-D (as well as Sec. III-C).
We examine the performance vs. FLOPs plot for the models, where in this plot the vertical axis is the performance of the model (e.g., RMSE for the pixel-wise regression EO task in Fig. 8, or accuracy for the semantic segmentation classification task in Fig. 9), and the horizontal axis is the number of FLOPs needed by the model for the downstream task. In Fig. 8, the results are for road and building density pixel-wise regression, and we also examine the performance and FLOPs for land cover semantic segmentation in Fig. 9. For the latter, we observe that for this specific task, the best models are TerraMind (i.e. see Sec. V-E) and Geo-Aware U-Net 44M.
For reference, at this point, we also provide the FLOPs for the usually used ResNet models (i.e. these are common classification models and not EO FMs): ResNet-18 1.191.19G FLOPs, 1212M parameters, ResNet-50 2.682.68G FLOPs, 2626M model parameters, and ResNet-152 7.547.54G FLOPs, 6060M parameters. We also note that the analysis of the computational complexity using FLOPs in Figs. 8 and 9 also indicates that sometimes smaller models can be better than larger models, i.e. a claim that, for example, has also recently been made in [64].
Performance of ViT vs. U-Net CNN. Taking into account all the three downstream tasks, i.e. Figs. 7-11, for $n$-shots larger than $1000$, i.e. for a large $n$-shot samples number, ViT UPerNet is better than the U-Net CNN model, and this is important. This is mainly because of the performance of the ViT on the semantic segmentation land cover mapping downstream task (i.e. the model ViT UPerNet 300M-0.5T, black colour in the figure vs. the model U-Net 200M-2T, green colour), as well as because of the performance of the ViT on the pixel-wise regression road and building density estimation.
@@
The effect of increasing the model size. In Figs. 7 and 10, we also examine the results when increasing the model size together with increasing the size of the pretraining dataset. This is for the 200M parameters model, the Geo-Aware U-Net 200M-2T model in the figures. EO FMs pretrained on general-purpose data, when scaled appropriately (i.e. joint increase of pretraining dataset size and model size), can eventually outperform models pretrained on specialized EO data.
Building density regression. As shown in Fig. 10, similar trends are observed for building density estimation. In particular, at $n=100$ and $500$, both the PhilEO 44M MajorTOM and FastTOM models outperform in RMSE the PhilEO Globe 44M baseline. In addition, the PhilEO 200M FastTOM model again performs best across most $n$-shot settings.
@@
Figure 12: Road density estimation in RMSE, comparing the models: the best ViT model with the Mamba models.
Figure 13: Building density in RMSE for the best ViT and the Mamba models.
Evaluation results, and land cover being correlated with seasonality. The dataset PhilEO Globe 0.5TB is multi-temporal and includes four time steps corresponding to the four seasons of the year. In Fig. 11, we observe much more correlation between land cover and including the four seasons of the year in the training set compared to between road density and including the four seasons in the training dataset, as well as between building density and including seasonality in the pretraining dataset, in Figs. 7 and Fig. 10 respectively.
The main question of which is the best model. In this work, we are also interested in which is the best model. This depends on the task (as well as on the $n$-shot samples number—e.g., see Fig. 7—), and more specifically on whether we focus on pixel-wise regression or semantic segmentation. For the former, the best model for road and building density estimation is the model Geo-Aware U-Net 200M (i.e. Figs. 7 and 10, as well as Fig. 8). In addition, for the latter, for the task of land cover mapping, TerraMind is the best model (see Sec. V-E, as well as Fig. 9), and the main reason for this is because in its modalities in the pretraining (i.e. in its correlation learning process), it includes land cover classes, geo-location, DEM, S-1 SAR, and NDVI data in addition to S-2 multi-spectral images. Here, the next best model for land cover classification is the model Geo-Aware U-Net 44M.
To better understand the impact of general-purpose versus specialized EO pretraining, we compare the performance of models trained on specialized land-only datasets (i.e. PhilEO Globe 0.5TB and FastTOM 2TB) with the performance of models trained on a large general-purpose dataset (MajorTOM 23TB) which includes oceans and ice. Given that all models share the same Geo-Aware U-Net architecture (approximately 44M parameters) and that the smaller datasets use a land mask to exclude oceans and ice, we expect models trained on PhilEO Globe and FastTOM to perform better in downstream land-focused tasks. Indeed, in most cases, this expectation holds. However, for the road density estimation task under low-shot (few-shot) learning settings, the model pretrained on MajorTOM surprisingly outperforms the specialized models in terms of RMSE. This suggests that despite the broader and potentially noisier pretraining data, large-scale general pretraining can be advantageous under certain conditions.
Figure 14: Land cover mapping results in accuracy for the different models, i.e. for the best ViT model and the Mamba models, at various $n$-shots.
This study also allows us to analyze two key factors:
  • Specialization effect: Does pretraining on land-only images give an advantage when the downstream tasks are also land-focused?